OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 3(2x − 1) greater than or equal to 4(2x − 3) − 3? x less than or greater to 6 6 greater than or equal to x A number line with a closed circle on 6 and shading to the left. A number line with a closed circle on 6 and shading to the right. best answer gets a medal! which one is the right answer??

OpenStudy (campbell_st):

have you attempted to solve it \[3(2x - 1)\ge4(2x - 3) - 3\] start by distributing \[6x - 3 \ge 8x - 12 - 3\] and continue from here...

OpenStudy (anonymous):

yeah ive tried a few times

OpenStudy (stefrheart):

What are you having trouble with?

OpenStudy (anonymous):

this problem..

OpenStudy (anonymous):

like is it less than or greater than i just dont know

OpenStudy (stefrheart):

would you like a step by step?

OpenStudy (anonymous):

yeah!

OpenStudy (stefrheart):

lol okay \[3\left( 2x-1 \right)\ge4\left( 2x-3 \right)-3\] \[6x-3\ge8x-12-3\] \[6x-3\ge8x-15\]

OpenStudy (stefrheart):

\[6x+12\ge8x\]

OpenStudy (stefrheart):

\[12\ge2x\]

OpenStudy (stefrheart):

\[6\ge x or x \le6\]

OpenStudy (anonymous):

thnx that rlly helped lol

OpenStudy (stefrheart):

welcome!